The nutrition information on the cereal box says that a 1/3 cup serving provides 140 calories and 6 grams of fiber. At that rate find how many calories and grams of fiber are in 1/2 cup serving..
Calories??
Fiber??

Found 2 solutions by stanbon, josmiceli:
Answer by stanbon(75887)

The nutrition information on the cereal box says that a 1/3 cup serving provides 140 calories and 6 grams of fiber.
At that rate find how many calories and grams of fiber are in 1/2 cup serving..
(1/3)x = 1/2
x = (1/2)/(1/3) = 3/2 = 1.5
---
There are 1.5 of those 1/3 cups in 1/2 cup.
----
Calories: 1.5(140) = 210 calories
Fiber: 1.5(6) = 9 grams of fiber
---------------------------
Cheers,
Stan H.




Answer by josmiceli(19441)

You can solve in steps
A 1/3 cup serving provides 140 calories and 6 grams of fiber.
A 1 cup serving would have 3 x 140 calories and 3 x 6 grams of fiber
That is, 420 calories and 18 grams fiber
A 1/2 cup would contain half of that, or
1/2 x 420 = 210 calories and 1/2 x 18 = 9 grams of fiber